# Contract Design for Pandora’s Box

### v2

@@ Abstract
We study a natural application of contract design to search problems with probabilistic prior and exploration costs. These problems have a plethora of applications and are expressed concisely within the Pandora’s Box model. Its optimal solution is the ingenious index policy proposed originally by Weitzman [39].
In our principal–agent setting, the search task is delegated to an agent. The agent performs a sequential exploration of $n$ boxes, suffers the exploration cost for each inspected box, and selects the content (called the prize) of one inspected box as outcome. Agent and principal obtain an individual value based on the selected prize. To influence the search, the principal a-priori designs a contract with a non-negative payment to the agent for each potential prize. The goal of the principal to maximize her expected reward, i.e., value minus payment. We show how to compute optimal contracts for the principal in several scenarios.
A popular and important subclass are linear contracts, and we show how to compute optimal linear contracts in polynomial time. For general contracts, we consider the standard assumption that the agent suffers cost but obtains value only from the transfers by the principal. Interestingly, a suitable adaptation of the index policy results in an optimal contract here. More generally, for general contracts with non-zero agent values for outcomes we show how to compute an optimal contract in two cases: (1) when each box has only one prize with non-zero value for principal and agent, (2) for i.i.d. boxes with a single prize with positive value for the principal. These results show that optimal contracts can be highly non-trivial, and their design goes significantly beyond the application or re-interpretation of the index policy.
@@ 1 Introduction
In many real-world situations, e.g., the search for a job candidate or house, a decision maker is faced with the choice between different alternatives of a-priori unknown value, which can be explored at some cost. A large class of such settings can be modeled by the famous Pandora’s box model proposed and analyzed by Weitzman [39]. In this model, there are $n$ alternatives or boxes. The value in each box is distributed independently according to a known probability distribution. There is a known additive function for the exploration cost of the boxes. The goal is to maximize the value of the selected box minus the cost of explored boxes. Starting from [33], this problem has recently received a lot of attention at the intersection of economics and computer science (see, e.g., [38, 26, 8, 10, 6, 25] and a recent survey [7]).
Due to missing qualifications or time constraints, in many markets the aforementioned exploration tasks are not executed directly by the decision maker. Rather, a decision maker (called principal $\mathcal{P}$ in the following) can delegate the exploration to an agent $\mathcal{A}$, whose incentives are potentially misaligned with that of $\mathcal{P}$.
@@
In this work, we initiate the study of algorithmic contract design for the Pandora’s Box problem. Specifically, we consider the following problem (for a fully formal definition, see Section 2): There are $n$ boxes, the $i$-th of which contains a prize, drawn independently from a known probability distribution, and can be opened at known cost $c_{i}$. The $j$-th prize from box $i$ has some utilities $a_{ij}$ and $b_{ij}$ for $\mathcal{A}$ and $\mathcal{P}$, respectively. Based on that information, $\mathcal{P}$ commits to a contract, specifying a transfer $t_{ij}\geq 0$ for each prize, where we assume non-negative transfers (limited liability). Given such a contract, $\mathcal{A}$ faces an instance of the classic Pandora’s Box Problem, which they solve optimally. Assuming $\mathcal{A}$ selects prize $j$ from box $i$, the resulting final utility of $\mathcal{P}$ is $b_{ij}-t_{ij}$. The goal is to efficiently compute a contract that maximizes the expectation of the latter quantity. Notably, ties in the optimization of $\mathcal{A}$ are broken in favor of $\mathcal{P}$, which is a standard assumption that makes the maximization problem well-defined.
Optimal contracts can be computed in time polynomial in the number of outcomes and the number of actions of the agent, using linear programming (cf., e.g., [20]). Note that this result does not yield a polynomial-time approximation algorithm for our problem since the action space is extremely succinctly represented—its size is lower-bounded by the number of opening orders, $n!$.
Nevertheless, we show that in a variety of standard cases, optimal contracts can be computed in polynomial time (such as, e.g., optimal linear contracts, or optimal contracts when the agent has no intrinsic value). Our results leave open the question whether computing optimal contracts efficiently is always possible in general – but our insights suggest that doing so is a highly non-trivial task and represents one of several very interesting avenues for future research.
@@
Our results build on Weitzman’s characterization of optimal policies for the original Pandora’s Box problem [39]. We reformulate it for $\mathcal{A}$’s subproblem: The fair cap of box $i$ is a value such that the expected amount of $a_{ij}+t_{ij}$ exceeding the fair cap is precisely $c_{i}$. The boxes are considered in non-increasing order of fair caps. If the best $a_{ij}+t_{ij}$ (initially 0) observed so far is larger than the next fair cap, the next box is opened; if it is smaller, a corresponding prize is accepted. If neither is the case, $\mathcal{A}$ is indifferent.
@@
The first observation we make is that Weitzman’s policy does actually not solve $\mathcal{A}$’s problem entirely. The reason is that we assume that $\mathcal{A}$ breaks ties in favor of $\mathcal{P}$. Note that the number of boxes that share some given fair cap $\varphi$ may be linear in $n$, leading to a number of possible ways of choosing the order that is exponential in $n$. Indeed, the order of such boxes may matter to $\mathcal{P}$ since different orders may lead to stopping with prizes that lead to vastly different utilities for $\mathcal{P}$. A complicating factor is that ties whether to select a prize with agent utility equal to $\varphi$ have to be broken simultaneously. Interestingly, we show that the problem of breaking ties can be reduced (in polynomial time) to solving a different instance of the classic Pandora’s box problem.
@@
With a solution of $\mathcal{A}$’s problem at hand, the main trade-off inherent to an optimal contract can be understood as follows: On the one hand, transfers manipulate the fair caps, the resulting exploration order, and the acceptance decisions of $\mathcal{A}$ in such a way that leads to a more favorable outcome for $\mathcal{P}$. On the other hand, transfers make $\mathcal{P}$ lose utility since they are costly.
@@
We start by considering linear contracts [30, 20], a class of contracts that has been particularly successful due to its simplicity. Here, a contract is fully described by a number $\alpha\in[0,1]$, which defines transfers $t_{ij}=\alpha\cdot b_{ij}$. We identify a polynomial-time computable set of critical values of $\alpha$ at which the set of $\mathcal{A}$-optimal policies (i.e., the set of policies that satisfy the above description) changes. By reasoning about the behavior of $\mathcal{P}$’s utility between any two such values, we can argue that the optimal contract needs to be a critical value. This type of argument is reminiscent of arguments from the recent literature (e.g., [17, 22, 16]), but here it requires special care. In particular, due to the tie breaking in favor of $\mathcal{P}$, which—as discussed above—is very relevant here, we need to keep track of the set of $\mathcal{A}$-optimal policies.
@@
Moving on to general contracts, we first consider the standard assumption that $a_{ij}=0$, i.e., that the agent has no intrinsic motivation to explore any boxes and is only motivated by payments from $\mathcal{P}$. We can again reduce the problem to the original Pandora’s Box problem! Observe that the utility that $\mathcal{P}$ extracts by solving the problem themselves (i.e., ignoring the agent and paying the cost themselves) is a straightforward baseline; no policy can extract more value for $\mathcal{P}$. Solving problem from $\mathcal{P}$’s perspective yields fair caps $\smash{\varphi_{i}^{\mathcal{P}}}$. It is not difficult to show that, by transferring any value exceeding $\smash{\varphi_{i}^{\mathcal{P}}}$ to $\mathcal{A}$, the $\mathcal{P}$-optimal policy can be enforced without a loss in utility for $\mathcal{P}$. The idea is somewhat inspired by [33] for the original Pandora’s Box problem but different.
@@
An orthogonal special case that has been considered in the literature [5] is that of binary boxes. That is, each box contains one of only two prizes, and one of them has 00 value for both $\mathcal{P}$ and $\mathcal{A}$ while the other one is arbitrary. Without any payments, the fair cap (from $\mathcal{A}$’s perspective again) of each box is its basic fair cap. The principal now faces the question of whether to lift (by transfers) the basic fair caps of favorable boxes above (or up to) the fair caps of less favorable boxes. While the latter is true in general, the special structure of binary boxes allows us to consider boxes in non-increasing order of their basic fair caps and move their fair cap greedily. Note that this algorithm alone does not suffice to tackle more general cases. Indeed, when boxes are not binary, it may be profitable for the principal to choose different payments for two boxes with the same distribution and the same fair caps – a condition we encounter in the next and final case.
@@
The final case we consider has a more general value structure, but we compromise on the boxes having different distributions. Specifically, we consider instances with only single positive prize for $\mathcal{P}$, and all distributions are identical. Still, transfers can be made depending on the box from which the prize originates. Again, the complexity is reduced by the fact that only a single payment per box has to be decided. Through an intricate sequence of exchange arguments and the help of continuization, we establish that there is an optimal contract with a simple structure: (i) In a first phase, the prize with positive principal value gets immediately accepted, and all transfers are identical; (ii) all transfers for boxes in the second phase are also identical. These conditions allow to find the optimal contract by enumerating a polynomial-sized set of contracts.
Generalizing our results and identifying hard classes of instances, if they exist, is an exciting direction for future research.
@@
After a review of related work in the subsequent section, we formally introduce our problem in Section 2 along with preliminary observations. Section 3 treats linear contracts. Section 4 deals with general contracts in several special cases (no agent value in Section 4.1 and binary boxes in Section 4.2). For spatial reasons, general contracts for i.i.d. boxes with single positive prize are discussed in Appendix C. Missing proofs are provided in Appendices A and B.
@@ 1.1 Related Work
Delegation is a related approach in the principal–agent framework and also received a lot of attention in the economics literature starting from seminal work of Holstrom [31]. Computational aspects of delegation are receiving interest recently, initiated by Kleinberg and Kleinberg [32], In these models, the principal $\mathcal{P}$ delegates a (search) task to an agent $\mathcal{A}$, again with potentially different interests, who has to inspect alternatives and propose an observed prize to $\mathcal{P}$. Instead of committing to a contract with payments, $\mathcal{P}$ limits the set of prizes she is willing to accept upon proposal by $\mathcal{A}$. A prominent objective is to find good acceptance policy for $\mathcal{P}$ and bound their performance by the delegation gap, which measures the multiplicative loss in utility for $\mathcal{P}$ in comparison to the case when $\mathcal{P}$ performs the (undelegated) search problem themselves. On a technical level, there are close connections to prophet inequalities [32] and online contention resolution schemes, which were established in different variants, such as multiple agents [37], stochastic probing [4, 5], as an online problem [11], or with limited information about agent utilities [13].
@@ 3 Optimal Linear Contracts
We now call $\alpha\in[0,1]$ a critical value if $\alpha\in\{0,1\}$ or one of the following properties is fulfilled:
@@
    There exist and $i,i^{\prime}\in[n]$ with $c_{i}>0,c_{i^{\prime}}>0$ and $j\in[m]$ such that the order between the value of prize $(i,j)$ for $\mathcal{A}$ and the fair cap of box $i^{\prime}$ changes. Formally, $v_{ij}(\alpha)=\smash{\varphi_{i^{\prime}}}(\alpha)$, and there exists $\varepsilon>0$ such that $v_{ij}(\alpha^{\prime})\neq\smash{\varphi_{i^{\prime}}}(\alpha^{\prime})$ for all $\alpha^{\prime}\in(\alpha-\varepsilon,\alpha)$ or for all $\alpha^{\prime}\in(\alpha,\alpha+\varepsilon)$.
@@
According to Lemma 1, there are at most $\mathcal{O}(m)$ critical values induced by case (c). By the same lemma, there are at most $\mathcal{O}(m)$ critical values for every box $i\in[n]$ with $c_{i}>0$ induced by case (b), as $\smash{\varphi_{i}}$ is convex and there are $m$ affine functions $v_{i^{\prime}j}$. Thus, $\mathcal{O}(nm)$ critical values in total are due to cases (b) and (c). Similarly, each linear segment of $\smash{\varphi_{i}}$ can intersect with another convex function $\smash{\varphi_{i^{\prime}}}$ at most twice. As there are at $\mathcal{O}(m)$ linear segments for $\alpha\in[0,1]$ and $n$ possible functions $\smash{\varphi_{i^{\prime}}}$, there are at most $\mathcal{O}(nm)$ such intersections for $\smash{\varphi_{i}}$. Thus, case (a) induces at most $\mathcal{O}(n^{2}m)$ critical values. Overall, the number of critical values is polynomial in $n$ and $m$, and the set of them can be computed in polynomial time.

### v3

# Designing Exploration Contracts

@@ Abstract
We study a natural application of contract design in the context of sequential exploration problems. In our principal–agent setting, a search task is delegated to an agent. The agent performs a sequential exploration of $n$ boxes, suffers the exploration cost for each inspected box, and selects the content (called the prize) of one inspected box as outcome. Agent and principal obtain an individual value based on the selected prize. To influence the search, the principal a-priori designs a contract with a non-negative payment to the agent for each potential prize. The goal of the principal is to maximize her expected reward, i.e., value minus payment. Interestingly, this natural contract scenario shares close relations with the Pandora’s Box problem.
We show how to compute optimal contracts for the principal in several scenarios. A popular and important subclass is that of linear contracts, and we show how to compute optimal linear contracts in polynomial time. For general contracts, we obtain optimal contracts under the standard assumption that the agent suffers cost but obtains value only from the transfers by the principal. More generally, for general contracts with non-zero agent values for outcomes we show how to compute an optimal contract in two cases: (1) when each box has only one prize with non-zero value for principal and agent, (2) for i.i.d. boxes with a single prize with positive value for the principal.
@@ 1 Introduction
In many real-world situations, e.g., the search for a job candidate or house, a decision maker is faced with the choice between different alternatives of a-priori unknown value, which can be explored at some cost. Due to missing qualifications or time constraints, in many markets such exploration tasks are not executed directly by the decision maker. Rather, a decision maker (called principal $\mathcal{P}$ in the following) can delegate the exploration to an agent $\mathcal{A}$, whose incentives are potentially misaligned with that of $\mathcal{P}$.
@@
In this work, we initiate the study of algorithmic contract design for exploration problems. Specifically, we consider the following problem (for a fully formal definition, see Section 2): There are $n$ boxes, the $i$-th of which contains a prize, drawn independently from a known probability distribution, and can be opened by paying a known exploration cost $c_{i}$. The $j$-th prize from box $i$ has some utilities $a_{ij}$ and $b_{ij}$ for $\mathcal{A}$ and $\mathcal{P}$, respectively. Based on that information, $\mathcal{P}$ commits to a contract, specifying a transfer $t_{ij}$ for each prize, where we assume non-negative transfers (limited liability). Given such a contract, $\mathcal{A}$ picks an adaptive search strategy that allows to optimize the trade-off between value of the selected prize (for $\mathcal{A}$) and total exploration cost. $\mathcal{A}$ uses the strategy to sequentially open boxes by paying the exploration cost. Depending on the observed prize, it determines the next box to open or to stop by selecting a prize from an open box. Assuming $\mathcal{A}$ selects prize $j$ from box $i$, the resulting final utility of $\mathcal{P}$ is $b_{ij}-t_{ij}$. The goal is to efficiently compute a contract that maximizes the expectation of the latter quantity. Here we assume that ties in the optimization of $\mathcal{A}$ are broken in favor of $\mathcal{P}$, which is a standard assumption that makes the maximization problem well-defined.
For the problem of computing optimal contracts, there is a solution by linear programming that runs in time polynomial in the number of outcomes and the number of actions of the agent (cf., e.g., [20]). Note that this result does not yield a polynomial-time algorithm for our problem since the action space is extremely succinctly represented—its size is lower-bounded by the number of opening orders, $n!$.
@@
We show that, in a variety of natural cases, optimal contracts can be computed in polynomial time (such as, e.g., optimal linear contracts, or optimal contracts when the agent has no intrinsic value). While our results leave open the question whether computing optimal contracts efficiently is always possible in general, our insights suggest that doing so is a highly non-trivial task and represents one (of several) very interesting avenue for future research.
@@
Notably, $\mathcal{A}$ faces an exploration problem known in the literature as the Pandora’s Box Problem. It was famously proposed and analyzed by Weitzman [39], who characterized optimal search strategies by a simple greedy rule. Starting from [33], this problem has recently received a lot of attention at the intersection of economics and computer science (see, e.g., [38, 26, 8, 10, 6, 25] and a recent survey [7]). We reformulate Weitzman’s characterization for $\mathcal{A}$’s task: The fair cap of box $i$ is a value such that the expected amount of $a_{ij}+t_{ij}$ exceeding the fair cap is precisely $c_{i}$. The boxes are considered in non-increasing order of fair caps. If the best $a_{ij}+t_{ij}$ (initially 0) observed so far is larger than the next fair cap, the next box is opened; if it is smaller, a corresponding prize is accepted. If neither is the case, $\mathcal{A}$ is indifferent.
@@
The first observation we make is that Weitzman’s policy does not solve $\mathcal{A}$’s problem entirely. The reason is that $\mathcal{A}$ breaks ties in favor of $\mathcal{P}$—recall that this is the standard assumption that makes the maximization problem well-defined. Note that the number of boxes that share some given fair cap $\varphi$ may be linear in $n$, leading to a number of possible ways of choosing the order that is exponential in $n$. Indeed, the order of such boxes may matter to $\mathcal{P}$ since different orders may lead to stopping with prizes that lead to vastly different utilities for $\mathcal{P}$. A complicating factor is that ties whether to select a prize with agent utility equal to $\varphi$ have to be broken simultaneously. Interestingly, we show that the problem of breaking ties can be reduced—in polynomial time but in a non-obvious way—to solving different instances of the classic Pandora’s box problem.
@@
With a solution of $\mathcal{A}$’s problem at hand, the main trade-off inherent to an optimal contract is as follows: Transfers manipulate the fair caps, the resulting exploration order, and the acceptance decisions of $\mathcal{A}$ in such a way that leads to a more favorable outcome for $\mathcal{P}$ – but transfers are costly for $\mathcal{P}$ which is unfavorable.
We start by considering linear contracts [30, 20], a class of simple and very intuitive contracts that has received a lot of attention due to its widespread use in applications. The agent receives a constant fraction of the principal’s revenue as commission. Formally, a linear contract is represented by a number $\alpha\in[0,1]$, which defines transfers $t_{ij}=\alpha\cdot b_{ij}$.
@@
To compute linear exploration contracts, we identify a polynomial-time computable set of critical values of $\alpha$ at which the set of $\mathcal{A}$-optimal policies (i.e., the set of policies that satisfy the above description) changes. By reasoning about the behavior of $\mathcal{P}$’s utility between any two such values, we can argue that the optimal contract needs to be a critical value. This type of argument is reminiscent of arguments from the recent literature (e.g., [17, 22, 16]), but here it requires special care. Due to the tie breaking in favor of $\mathcal{P}$, which—as discussed above—is a key property here, we need to keep track of the set of $\mathcal{A}$-optimal policies.
For standard contract design, the authors in [20] describe a class of instances in which the restriction to linear contracts leads to a multiplicative loss of $n$ in $\mathcal{P}$’s achievable utility, where $n$ is the number of actions. In this class of instances, action $i$ with cost $c_{i}$ deterministically leads to an individual outcome with value $R_{i}$. We can create one box for each action $i$, with opening cost $c_{i}$ and a single outcome with agent utility 00 and principal utility $R_{i}$. This leads to a straightforward one-to-one correspondence between contracts in the two instances that preserves utilities and linearity. Thus, the multiplicative loss of $n$ transfers to our model, which motivates our search for optimal general exploration contracts.
We first consider general contracts under the standard assumption that $a_{ij}=0$, i.e., that the agent has no intrinsic motivation to explore any boxes and is only motivated by payments from $\mathcal{P}$. In contract design, this assumption is usually without loss of generality, since arbitrary intrinsic agent values $a_{ij}\neq 0$ can be accounted for with action costs to obtain an equivalent instance with agent values $a_{ij}=0$. However, for the exploration contracts we consider, such an adjustment changes the costs of exploration strategies (rather than individual boxes). This substantially changes the structure of the problem. Indeed, as we see below, optimal policies for the problem with non-zero agent value have substantially different properties than the ones for zero value.
@@
Clearly, the utility that $\mathcal{P}$ can extract by solving the exploration problem themselves (i.e., ignoring the agent and paying the cost themselves) is a straightforward baseline; no policy can extract more value for $\mathcal{P}$. Solving the problem from $\mathcal{P}$’s perspective yields fair caps $\smash{\varphi_{i}^{\mathcal{P}}}$. If all $a_{ij}=0$, then by transferring any value exceeding $\smash{\varphi_{i}^{\mathcal{P}}}$ to $\mathcal{A}$, the $\mathcal{P}$-optimal policy is adopted by $\mathcal{A}$, without a loss in utility for $\mathcal{P}$. This implies, in particular, that delegating the exploration to $\mathcal{A}$ using an optimal contract yields the same utility for $\mathcal{P}$ as if she performed the exploration on her own. Formally, the idea is somewhat inspired by [33] for the original Pandora’s Box problem but different.
@@
An orthogonal special case that has been considered in the literature (see, e.g., [5]) is that of binary boxes. That is, each box contains one of only two prizes, and one of them has 00 value for both $\mathcal{P}$ and $\mathcal{A}$ while the other one is arbitrary. Without any payments, the fair cap (from $\mathcal{A}$’s perspective again) of each box is its basic fair cap. The principal now faces the question of whether to lift (by transfers) the basic fair caps of favorable boxes above (or up to) the fair caps of less favorable boxes. While the latter is true in general, the special structure of binary boxes allows us to compute an optimal exploration contract by considering boxes in non-increasing order of their basic fair caps and moving their fair cap greedily. Note that this algorithm alone does not suffice to tackle more general cases. Indeed, when boxes are not binary, it may be profitable for the principal to choose different payments for two boxes with the same distribution and the same fair caps – a condition we encounter in the next and final case.
@@
The final case we consider has a more general value structure, but we compromise on the boxes having different distributions. Specifically, we consider instances with only a single positive prize for $\mathcal{P}$, and all distributions are identical. Still, transfers can be made depending on the box from which the prize originates. Again, the complexity is reduced by the fact that only a single payment per box has to be decided. Through an intricate sequence of exchange arguments and the help of continuization, we establish that there is an optimal contract with a simple structure: (i) In a first phase, the prize with positive principal value gets immediately accepted, and all transfers are identical; (ii) all transfers for boxes in the second phase are also identical. These conditions allow to find the optimal contract by enumerating a polynomial-sized set of contracts.
In this final case, we also give a family of instances in which both the above phases exist and have substantial lengths in any optimal solution. Hence, perhaps surprisingly, such a relatively simple case already has optimal contracts with a fairly complicated structure.
@@
After a review of related work in the subsequent section, we formally introduce our problem in Section 2 along with preliminary observations. Section 3 treats linear contracts. Section 4 deals with general contracts in several special cases (no agent value in Section 4.1 and binary boxes in Section 4.2). General contracts for i.i.d. boxes with a single positive prize are discussed in Section 4.3. Missing proofs are provided in Appendices AB and C.
@@ 1.1 Related Work
Delegation is a related approach in the principal–agent framework and also received a lot of attention in the economics literature starting from seminal work of Holstrom [31]. Computational aspects of delegation are receiving interest recently, initiated by Kleinberg and Kleinberg [32]. In these models, the principal $\mathcal{P}$ delegates a (search) task to an agent $\mathcal{A}$, again with potentially different interests, who has to inspect alternatives and propose an observed prize to $\mathcal{P}$. Instead of committing to a contract with payments, $\mathcal{P}$ limits the set of prizes she is willing to accept upon proposal by $\mathcal{A}$. A prominent objective is to find good acceptance policy for $\mathcal{P}$ and bound their performance by the delegation gap, which measures the multiplicative loss in utility for $\mathcal{P}$ in comparison to the case when $\mathcal{P}$ performs the (undelegated) search problem themselves. On a technical level, there are close connections to prophet inequalities [32] and online contention resolution schemes, which were established in different variants, such as multiple agents [37], stochastic probing [4, 5], as an online problem [11], or with limited information about agent utilities [13]. Perhaps most related is a version of the problem studied in [5]. Here, a principal delegates an instance of the Pandora’s Box problem by committing to a set of acceptable outcomes. An agent has to perform the costly inspection of alternatives. This standard delegation setting has some obvious limitations, e.g., the agent would never perform an inspection if his expected intrinsic value does not cover the inspection cost. A model variant with transfers is also studied in [5] in which the principal can reimburse the agent for exploration costs. Crucially, this is impossible in hidden-action principal-agent settings that we consider, where $\mathcal{P}$ simply cannot directly reimburse $\mathcal{A}$ for his expenses, since she cannot observe which inspections have been performed by $\mathcal{A}$.
@@ 3 Optimal Linear Contracts
We call $\alpha\in[0,1]$ a critical value if $\alpha\in\{0,1\}$ or one of the following properties holds:
@@
    There exist $i,i^{\prime}\in[n]$ with $c_{i}>0,c_{i^{\prime}}>0$ and $j\in[m]$ such that the order between the value of prize $(i,j)$ for $\mathcal{A}$ and the fair cap of box $i^{\prime}$ changes. Formally, $v_{ij}(\alpha)=\smash{\varphi_{i^{\prime}}}(\alpha)$, and there exists $\varepsilon>0$ such that $v_{ij}(\alpha^{\prime})\neq\smash{\varphi_{i^{\prime}}}(\alpha^{\prime})$ for all $\alpha^{\prime}\in(\alpha-\varepsilon,\alpha)$ or for all $\alpha^{\prime}\in(\alpha,\alpha+\varepsilon)$.
@@
According to Lemma 1, there are at most $\mathcal{O}(n)$ critical values induced by case (c). By the same lemma, there are at most $\mathcal{O}(nm)$ critical values for every box $i\in[n]$ with $c_{i}>0$ induced by case (b), as $\smash{\varphi_{i}}$ is convex and there are $nm$ affine functions $v_{i^{\prime}j}$, each of which intersects $\smash{\varphi_{i}}$ at most twice. Thus, $\mathcal{O}(n^{2}m)$ critical values in total are due to cases (b) and (c). Similarly, each linear segment of $\smash{\varphi_{i}}$ can intersect with another convex function $\smash{\varphi_{i^{\prime}}}$ at most twice. As there are at $\mathcal{O}(m)$ linear segments for $\alpha\in[0,1]$ and $n$ possible functions $\smash{\varphi_{i^{\prime}}}$, there are at most $\mathcal{O}(nm)$ such intersections for $\smash{\varphi_{i}}$. Thus, case (a) induces at most $\mathcal{O}(n^{2}m)$ critical values. Overall, the number of critical values is polynomial in $n$ and $m$, and the set of them can be computed in polynomial time.